OpenStudy (anonymous):

Let f(x) be the function 11x^2-6x+4. Then the quotient (f(10+h)-f(10))/h can be simplified to ah+b for: a=? b=? My textbook is really confusing and my professor isn't the best and I have a bunch of problems like these that I'm having a lot of trouble even getting started. If you could help it would be greatly appreciated.

OpenStudy (anonymous):

which is wrong but my best effort

OpenStudy (campbell_st):

ok... so you know \[f(x) = 11x^2 - 6x + 4\] so \[f(x + h) = 11(x + h)^2 - 6(x + h) + 4\] you need to now simplify this \[f(x + h) = 11(x^2 + 2xh + h^2) - 6x - 6h + 4\] which becomes \[f(x + h) = 11x^2 + 22xh + 11h^2 - 6x - 6h + 4\] so your problem now becomes \[\frac{f(x + h) - f(x)}{h} = \frac{11x^2 + 22xh + 4 -6x - 6h + 4 - 11x^2 + 6x - 4} {h}\] now you can simplify this by collecting like terms... hope it helps

OpenStudy (campbell_st):

oops should read, slight typo \[\frac{f(x + h) - f(x)}{h} = \frac{11x^2 + 22xh + h^2 - 6x - 6h + 4 - 11x^2 + 6x - 4}{h}\]

OpenStudy (anonymous):

Thank you, quick question. why isn't the h^2 11h^2?

OpenStudy (campbell_st):

ok... well remember you will divide by h so the will reduce... to jusy h... and you want ah + b so the simplified version is \[\frac{f(x + h) - f(x)}{h} = \frac{22xh +h^2 -6h}{h}\] so divide every term by h, and substitute x = 10

OpenStudy (anonymous):

=10

OpenStudy (campbell_st):

yes, the question says f(10 + h) so replace x with 10 and then simplify